# Enhanced Direction-Sensing Methods and Performance Analysis in Low-Altitude Wireless Network via a Rotation Antenna Array

### 21 Mar 2026 v1

@@ I System Model
II Derived CRLB for Rotatable Array
@@
By substituting Eq. (29) into Eq. (II), the elements of the FIM are reformulated as
@@
III Proposed Low-Complexity Rotatable Array DOA Methods
@@
III-A Proposed Low-Complexity PRI Method
@@
III-B Proposed Low-Complexity IGSS Method
@@
III-C Computational Complexity Analysis
@@ IV Simulation Results
Fig. 5-7 illustrate the RMSE convergence curves of the proposed low-complexity PRI and PRI-IGR methods for three representative incident angles, i.e., $\theta=5^{\circ},85^{\circ},175^{\circ}$. The CRLB is also included as the theoretical performance benchmark. It is seen that the CRLB is continuously decreased over iterations due to the accumulation of sampling signals. While acceptable at $\theta=5^{\circ}$, the conventional FA-Root-MUSIC method is completely disabled at large angles, i.e., $\theta=85^{\circ},175^{\circ}$, due to severe SNR attenuation. Although the RR-Root-MUSIC eventually converges, it requires multiple rotation iterations for larger boresight deflection angle ($\theta=85^{\circ},175^{\circ}$), resulting in high computational complexity, and still has a significant gap compared to the CRLB. By contrast, the proposed PRI method effectively improves the initialization accuracy by pre-rotation. On this basis, the PRI-IGSS method further improves the estimation through iterative greedy search with a continuous sampled signal accumulation process, leading to rapid convergence and performance that approaches the CRLB.
@@
Fig. 8 shows the performance curves of RMSE versus SNR of the proposed methods with $\theta=85^{o}$. Due to the directive antenna characteristics, the performance of the FA-Root-MUSIC is highly degraded. Although stable estimation is provided by the proposed PRI and traditional RR-Root-MUSIC respectively, and a noticeable gap from the CRLB is maintained. In contrast, the CRLB is perfectly achieved by the proposed PRI-IGSS across the entire SNR range from -15 dB to 35 dB. Consequently, one-order magnitude improvement in sensing accuracy is successfully achieved by PRI-IGSS compared with the PRI and RR-Root-MUSIC.
@@
In Fig. 9, the RMSE performance versus the elevation angle $\theta$ at a low SNR of -5 dB is evaluated. As $\theta$ increases or the target deviates from the array boresight, a severe performance degradation is suffered by the FA-Root-MUSIC, particularly in off-boresight regions, i.e., $\theta>80^{o}$. In contrast, the PRI and PRI-IGSS improve the estimation accuracy by about three and four orders of magnitude, respectively. Moreover, over the entire angular range, the PRI-IGSS consistently outperforms the RR-Root-MUSIC by about one order of magnitude. These results show that PRI-IGSS still exhibits excellent spatial stability and excellent estimation performance even in extreme cases.
Fig. 10 plots the RMSE performance versus the elevation angle $\theta$ for the different number of candidate pre-rotation directions at a SNR of -5 dB is evaluated. Based on equi-spaced rotation or sampling, the candidate angle sets for $Q=1,2,3,$ and $4$ are respectively defined as $\{90^{\circ}\}$, $\{45^{\circ},135^{\circ}\}$, $\{30^{\circ},90^{\circ},150^{\circ}\}$, and $\{22.5^{\circ},67.5^{\circ},112.5^{\circ},157.5^{\circ}\}$. As illustrated, a negligible performance difference between $Q=3$ and $Q=4$ is observed. This implies that the main high-gain region of the array is sufficiently covered by the candidate set of $Q=3$. Consequently, only marginal accuracy improvements are provided by further increasing $Q$, while the number of rotation is increased, i.e., from 9 to 16. Thus, $Q=3$ is verified as the optimal configuration to balance sensing precision and computational burden.
V Conclusions

### 03 Apr 2026 v3

@@ II System Model
III Derived CRLB for Rotatable Array
@@
By substituting Eq. (29) into Eq. (III), the elements of the FIM are reformulated as
@@
IV Proposed Low-Complexity Rotatable Array DOA Methods
@@
IV-A Proposed Low-Complexity PRI Method
@@
IV-B Proposed Low-Complexity IGSS Method
@@
IV-C Computational Complexity Analysis
@@ V Simulation Results
Fig. 5-7 illustrate the RMSE convergence curves of the proposed low-complexity PRI and PRI-IGR methods for three representative incident angles, i.e., $\theta=5^{\circ},85^{\circ},175^{\circ}$. The CRLB is also used as a performance benchmark. It is seen that the CRLB decreases as iteration process due to the accumulation of sampling signals. In Fig.5 with $\theta=5^{\circ}$, the conventional FA-Root-MUSIC method is completely disabled at large angles, i.e., $\theta=85^{\circ},175^{\circ}$, due to severe SNR induced by the directive property of array element. Although the RR-Root-MUSIC eventually converges, it requires multiple rotation iterations for larger boresight deflection angle ($\theta=85^{\circ},175^{\circ}$), resulting in high computational complexity, and still has a large gap compared to the CRLB. By contrast, the proposed PRI method effectively improves the initialization accuracy by pre-rotation. On this basis, the PRI-IGSS method further improves the estimation through iterative greedy search with a continuous sampled signal accumulation process, leading to a rapid convergence and an optimal that approaches the CRLB.
@@
Fig. 8 shows the performance curves of RMSE versus SNR of the proposed methods with $\theta=85^{o}$. Due to the directive antenna characteristics, the performance of the FA-Root-MUSIC degrades seriously for all SNR values. Although stable estimation is provided by the proposed PRI and traditional RR-Root-MUSIC respectively, and a noticeable gap from the CRLB is maintained. In contrast, the CRLB is perfectly achieved by the proposed PRI-IGSS across the entire SNR range from -15 dB to 35 dB. Consequently, almost an one-order magnitude improvement in sensing accuracy is successfully achieved by the proposed PRI-IGSS over the PRI and RR-Root-MUSIC.
@@
In Fig. 9, the RMSE performance versus the elevation angle $\theta$ at a low SNR of -5 dB is evaluated. As $\theta$ increases or the target deviates from the array boresight, a severe performance degradation is suffered by the FA-Root-MUSIC, particularly in off-boresight regions, i.e., $\theta>80^{o}$. In contrast, the proposed PRI and PRI-IGSS improve the estimation accuracy by about three and four orders of magnitude, respectively. Moreover, over the entire angular range, the PRI-IGSS consistently outperforms the RR-Root-MUSIC by about one-order magnitude. These results show that PRI-IGSS still exhibits excellent spatial stability and excellent estimation performance even in extreme cases.
Fig. 10 plots the RMSE performance versus the elevation angle $\theta$ for the different number of candidate pre-rotation directions at a SNR of -5 dB. Based on equi-spaced rotation or sampling, the candidate angle sets for $Q=1,2,3,$ and $4$ are respectively defined as $\{90^{\circ}\}$, $\{45^{\circ},135^{\circ}\}$, $\{30^{\circ},90^{\circ},150^{\circ}\}$, and $\{22.5^{\circ},67.5^{\circ},112.5^{\circ},157.5^{\circ}\}$. As illustrated, a negligible performance difference between $Q=3$ and $Q=4$ is observed. This implies that the main high-gain region of the array is sufficiently covered by the candidate set of $Q=3$. Consequently, only trivial accuracy improvements are reaped by further increasing $Q$, while the number of rotation is increased, i.e., from 9 to 16. Thus, $Q=3$ may be chosen as a good value to balance sensing precision and computational burden.
VI Conclusions